Two very small 8.80g spheres, 17.0cm apart from center to center, are charged by adding equal numbers of electrons to each of them.

Part A
-Disregarding all other forces, how many electrons would you have to add to each sphere so that the two spheres will accelerate at 23.0g when released?

N = ______________ electrons

Part B
-Which way will they accelerate?



Could somebody give a concise approach to this problem? Thank you.
Answer #1
Given that
mass of each sphere is m = 8.80x 10^-3 kg
acceleration due to gravity g =9.80 m/s2
----------------------------------------------------
We have
      F = ma
But electrostatic force is
            F = kq^2/r^2
    kq^2/r^2 =  ma
Here r = 8.5x10^-2 m
So the charge on the sphere is
           q = √[mar^2/k]
              = √[(8.80x 10^-3 kg)(23(9.8 m/s2))(8.5x10^-2 m)^2/(9.0x10^9 N.m2/C2)]
              = 1.26x10^-6C
The number of electrons are
            n =q/e
               = (1.26x10^-6C)/(1.6 x10^-19 C)
               = 7.88x10^+12
b) two sphere are repul each other